# Weyl-Einsteinian-Cubic Gravity

### v1

@@ III Weyl-Einsteinian Cubic Gravity
Now that we have presented primary operations, let us now proceed with our focal purpose of concretizing the WECG. In the construction, it will primarily be expected from WECG to recover the main attributes of usual ECG (e.g., the propagating degrees of freedom around constant curvature backgrounds reducing to that of GR at the linear level) in the subregion of the broken phase of WECG. Additionally, the relative coupling constants will be selected as those in the strongly coupled extension of ECG [19] upgraded here by the dimensionless free parameters $\alpha_{i}$ (with $i=1\cdots 8$) to get all the possible configurations of relative couplings among the higher order terms in the model. [Note that assuming $\alpha_{i}$’s will also become useful as one studies the unitary (ghost and tachyon-free) parameter regions of WECG]. Referring to [70] for the details of Weyl’s local conformal invariance via real scalar and abelian gauge fields, one can write the WECG gravity action with the help of properly tuned weight for the compensating scalar field as follows
@@
$g_{\mu\nu}\rightarrow g^{{}^{\prime}}_{\mu\nu}=e^{2\lambda(x)}g_{\mu\nu},\hskip 19.91684pt\Phi\rightarrow\Phi^{{}^{\prime}}=e^{-\lambda(x)}\Phi$
@@
$\widehat{R}^{\mu}{{}_{\nu\rho\sigma}}(g_{\mu\nu},A_{\mu})$ $=R^{\mu}{{}_{\nu\rho\sigma}}+\delta^{\mu}{{}_{\nu}}F_{\rho\sigma}+2\delta^{\mu}_{[\sigma}\nabla_{\rho]}A_{\nu}+2g_{\nu[\rho}\nabla_{\sigma]}A^{\mu}+2A_{[\sigma}\delta_{\rho]}^{\mu}A_{\nu}+2g_{\nu[\sigma}A_{\rho]}A^{\mu}$
@@
$S_{\Phi}=\pi\int d^{n}x\sqrt{-g}\Big{(}{\cal{D}}_{\mu}\Phi{\cal{D}}^{\mu}\Phi+\nu\,\Phi^{\frac{2n}{n-2}}\Big{)},\quad S_{A^{\mu}}=\varepsilon\int d^{n}x\sqrt{-g}\,\Phi^{\frac{2(n-4)}{n-2}}F_{\mu\nu}F^{\mu\nu}.$ (6)
Observe that $\nu\geq 0$ is a dimensionless coupling constant ensuring the presence of a ground state, and $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the ordinary abelian field strength tensor. Notice also that, due to being conformally invariant in $3+1$ dimensions, the compensating scalar field factor in Maxwell action disappears here [70]. In fact, one could leave out Maxwell’s kinetic term initially since such a term would also emerge on its own from the nature of the model. But, we will keep assuming it for the sake of generalization since the coefficient can always be tuned to get canonically normalized Maxwell’s action. Observe that, for the sake of future studies (say, unitarity analysis, black hole solutions, quantum corrections, etc.), all the dimensionless coefficients in (3) are taken to be arbitrary since extra restrictions on the relative couplings may be enforced by those prospective studies. At this point, one shall notice that it also deserves to generalize WECG in (3) (as well as its higher dimensional extensions) with the ensuing Weyl’s extension of the dimensionally-extended Euler quadratic and cubic densities (namely, $\widehat{\chi}_{4}$ and $\widehat{\chi}_{6}$) associated with Weyl-invariant analogous of Gauss-Bonnet and Cubic Lovelock combinations [19]
@@
$\hskip 17.07182pt+3\widehat{R}_{abcd}\widehat{R}^{abcd}R+24\widehat{R}_{abcd}\widehat{R}^{ac}\widehat{R}^{bd}+16\widehat{R}_{a}{{}^{b}}\widehat{R}_{b}{{}^{c}}\widehat{R}_{c}{{}^{a}}-12\widehat{R}_{ab}\widehat{R}^{ab}\widehat{R}+\widehat{R}^{3},$ (7)
@@
where $\mathfrak{L}_{\mbox{ECG}}$ is the Lagrangian density of the strongly-coupled ECG [19] upgraded by the dimensionles couplings $\alpha_{i}$ which involves specific cubic curvature combinations only and is described as below
@@ IV.1 Metric Field Equations
during the variation of the action with respect to $\delta g^{\mu\nu}$, one can show that, posterior to a long process of calculations, the metric field equation up to a boundary term can be written in the compact form according to the certain expressions of scalar field terms as follows
@@
Here, $\mathscr{O}_{i}=\mathscr{O}_{i}(R,A,\Phi)$ where $i=1\cdots 9$ in general. As sequentially presented below, each $\mathscr{O}_{i}$ term further contains the other kind of terms compactly denoted as “$\Theta_{[i]j}A=\Theta^{[i]j}A$” where the first indices “$i$” in the square bracket stands for specific $\mathscr{O}_{i}$-term in (12) it belongs to, and the second indice “$j$” signifies the number of the sub-term for same $\mathscr{O}_{i}$. Also, the pure specific curvature combinations in $\mathscr{O}_{i}$-terms are compactly denoted by $\widehat{\Omega}_{(i)}$ and presented in (B), explicitly. To be more precise, let us first notice that the term $\mathscr{O}^{(1)}_{\mu\nu}$ reads as
@@
[…]
@@
wherein the related term $\widehat{\Omega}^{(5)}_{\mu}$ comprising specific curvature interactions only is presented in (42). Here, the explicit forms of coupling $\kappa_{1({Q_{5}})}$ and coefficient functionals $\Theta_{\alpha}^{[5]i}A_{\beta}$ (with $i=1\cdots 21)$ are respectively exhibited in (D.3) and (C.5).
@@
where the relevant bare curvature term $\widehat{\Omega}^{(7)}_{\mu\nu}$ and function $\Xi^{\alpha\beta}_{(5)}$ can be found in (44) and (10), manifestly. The constants $\zeta_{i}$’s, $\kappa_{3(5)}$ and $\kappa_{j(7)}$ (with $j=5,6,9$) are presented in (A), (D.2) and (D.5), respectively. Also, the different coefficient functionals $\Theta_{[1]1}^{\alpha}A^{\beta}$, $\Theta_{[5]3}^{\alpha}A^{\beta}$, $\Theta^{\alpha}_{[6]k}A^{\beta}$ (for $k=2,4,10$) and $\Theta^{\alpha}_{[7]l}A^{\beta}$ (where $l=1\cdots 41$) are given in (C.1), (C.5), (C.6) and (C.7) in succession.
@@ IV.2 Field Equations for Weyl’s Vector and Scalar Fields
Firstly, due to the highly complex higher-order nature of the WECG as was underlined, we shall similarly collect separately the outcoming terms in the gauge field equations mainly according to $\nabla_{\mu}\Phi^{-2}$ and $\Phi^{-2}$ even though they are more simpler in contrast to the metric field equations. To be more precise, by varying the action in (8) with respect to $\delta A^{\mu}$, one gets the gauge field equation in compact form as follows
@@
Here, the explicit forms of coefficient functionals $\mathscr{X}_{[A]i}$ (where $i=1\cdots 11$) are given in (E).
As for the scalar field equation, it is actually straightforward to show that the variation of WECG action in (8) with respect to $\delta\Phi$ turns out to be
@@
V Maximally Symmetric Spacetime As Background Vacuum Solution
@@
[…]
In this part, we will see that the general field equations reduce to relatively simpler expressions around constant curvature backgrounds (or at least when the vacuum solutions are those of maximally symmetric spacetimes) although they apparently are highly intricate for generic backgrounds due to various higher-order nonminimal interactions among the existing fields ($g_{\mu\nu},A_{\mu},\Phi$). To see that clearly, let us also pick $F_{\mu\nu}=0$ with the specific legitimate ansatz $A_{\mu}=0$ to avoid having any certain direction that would break the Lorentz-invariance of vacua and also freeze the scalar field to its vacuum expectation value as $\Phi=\left<\Phi_{vac}\right>=\sqrt{{\cal M}}$ in addition to (24). With these proper setups, the vector field equation (22) will be fulfilled immediately, while the metric and scalar field equations in (12) and (23) yield the identical vacuum field equation incorporating background cosmological constant and vacuum expectation value of the scalar field–which is also approval of the calculations–as below
@@
which fixes one of the $11$ free parameters in the generic action. Recall that $\nu>0$ for a viable ground state. As for the solutions of (25), note that the process is bilateral. That is, one may proceed by presupposing that the vacuum expectation value of the scalar field ${\cal M}$ is known and then find the cosmological constant $\Lambda$ or vice versa. For the first case, by setting coefficient of the kinetic term of the scalar field to its canonical value (i.e., $\pi=-1/2$) and also using Cardano’s method for the depressed cubic equation, one gets the well-defined first generic solution for the background cosmological constant as
@@
where $\Delta_{0({\cal M})}=16\alpha^{2}_{1}\Lambda^{2}$ and the term ${\cal C}_{{\cal M}}$ turns out to be
@@
where $\widehat{\xi}$ and $r$ are given in (32). Observe that the two roots for both cases are imaginary so we will not study them in detail since they will provide non-hermitian complex operators associated with non-physical observables. (See conclusion for the relevant discussions and possible future directions.)
@@
Here, $\varphi_{c}=\left<\Phi_{vac}\right>$ is the classical or vacuum expectation value of $\Phi$ and $\hat{\mu}$ is the renormalized mass scale. Due to the present log singularity of $V^{eff}_{one-loop}$, $\hat{\nu}(\mu)$ must be defined at the point where the renormalization scale is different than zero. It is apparent that, unlike the tree-level, one-loop radiative contributions recast the effective potential (and so the couplings) such that the minimum of the potential is moved to a point where $\varphi_{c}\neq 0$. This shifting of the minimum designates a nonzero dimensionful vacuum expectation value to the scalar field that spontaneously breaks the existing local conformal symmetry. As in the ordinary Coleman-Weinberg mechanism, here the perturbation approach will collapse for the higher values of $\nu(\hat{\mu})$. However, this will likely be resolved once the vector fields are also considered, which is also the case in the usual Coleman-Weinberg mechanism. Nevertheless, the symmetry is broken by the dimensionful vacuum expectation value of the scalar field emerging due to the radiative corrections at the one-loop level, which is within the scope of the current paper. Here, as one of the future directions, notice that one needs to perform the Renormalization Group flow to determine the critical points and hence explicit viable regions for the relative coupling constants at the one-loop level in WECG.
@@ VI Conclusions and Discussions
By upgrading the bare curvature terms to those composed of abelian gauge fields along with appropriately tuned weights of real scalar fields, we have constructed a Weyl-gauging of strongly coupled Einsteinian-Cubic gravity in Weyl’s geometry in $3+1$ dimensions, called Weyl-Einsteinian-Cubic gravity (WECG). The model is invariant under the so-called Weyl’s local conformal transformations and hence does not involve any dimensionful parameter. Due to the nature of construction, the WECG describes $g_{\mu\nu},A_{\mu}$ and $\Phi$ as the fundamental dynamical fields. We have also obtained the most general field equations associated with the existing propagating fields about any generic spacetime background. Accordingly, we have demonstrated that the ordinary ECG emerges as the lower energy limit of WECG in its local non-conformal-invariant vacua about (anti-) de Sitter and flat spacetimes where the vacuum expectation value is fixed to the Planck mass scale. Additionally, we have gotten the same cubic vacuum field equations in the maximally symmetric backgrounds (namely, (anti-) de Sitter and flat spaces) from the generic field equations for the metric and real scalar field where that of the gauge field is automatically fulfilled as expected. Ensuingly, we have explicitly obtained the desired real cosmological constant and vacuum expectation value of the scalar field and also determined the reality conditions on the values of existing relative couplings in WECG for the (anti-) de Sitter vacuum spacetime. Here, the natural existence of (anti)-de Sitter space spontaneously breaks Weyl’s local conformal symmetry which fixes all the dimensionful parameters. Moreover, we have observed that anti-de Sitter space is allowed to be a vacuum background solution of WECG, while de Sitter space is ruled out. As for the flat spaces, the local conformal invariance of the vacuum persists to be unbroken. Here, we have demonstrated that the essential dimensionful parameter arises at the renormalization scale at the one-loop level, and thus the symmetry is broken radiatively.
@@
Acknowledgements
We thank Daniel Grumiller, Mithat Unsal, Daniel Kabat, Anupam Mazumdar, Paul Brown, Heeseung Zoe, Ercan Kilicarslan and Masoud Latifinavid for useful discussions and suggestions. This article is dedicated to Roman Jackiw’s first death anniversary who was one of the world-leading scientists of the century and my former postdoc advisor.

### v2

@@ III.2 Weyl-Invariant Extension of Einsteinian Cubic Gravity
As for our main model of WECG, the relative coupling constants will be selected as those in the strongly coupled (Cosmological) extension of ECG upgraded here by the dimensionless free parameters $\alpha_{i}$ (with $i=1\cdots 8$) to get all the possible configurations of relative couplings among the higher order terms in the model. [Note that assuming $\alpha_{i}$’s will also become useful as one studies the unitary (ghost and tachyon-free) parameter regions of WECG]. Referring to [71, 75] for the details of Weyl’s local conformal invariance via real scalar and abelian gauge fields, one can write the WECG gravity action with the help of properly tuned weight for the compensating scalar field as follows
@@
$g_{\mu\nu}\rightarrow g^{{}^{\prime}}_{\mu\nu}=e^{2\lambda(x)}g_{\mu\nu},\qquad\Phi\rightarrow\Phi^{{}^{\prime}}=e^{-\lambda(x)}\Phi$
${\cal{D}}_{\mu}\Phi=\partial_{\mu}\Phi-A_{\mu}\Phi,\hskip 25.6073ptA_{\mu}\rightarrow A^{{}^{\prime}}_{\mu}=A_{\mu}-\partial_{\mu}\lambda(x),$ (6)
@@
$S_{\Phi}=\pi\int d^{n}x\sqrt{-g}\Big{(}{\cal{D}}_{\mu}\Phi{\cal{D}}^{\mu}\Phi+\nu\,\Phi^{\frac{2n}{n-2}}\Big{)},\quad S_{A^{\mu}}=\varepsilon\int d^{n}x\sqrt{-g}\,\Phi^{\frac{2(n-4)}{n-2}}F_{\mu\nu}F^{\mu\nu},$ (8)
where $\nu\geq 0$ is a dimensionless coupling constant ensuring the presence of a ground state, and $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the ordinary abelian field strength tensor. Of course, one can also set the parameters $\pi$ and $\epsilon$ to their so-called canonically normalized values, but we shall keep them as free for now for the sake of future works such as the (tree-level) unitarity analysis. Notice also that, due to being conformally invariant in $3+1$ dimensions, the compensating scalar field factor in Maxwell action disappears here [71]. In fact, one could leave out Maxwell’s kinetic term initially since such a term would also emerge on its own from the nature of the model. But, we will keep assuming it for the sake of generalization since the coefficient can always be tuned to get canonically normalized Maxwell’s action. Observe that, for the sake of future studies (say, unitarity analysis, black hole solutions, quantum corrections, etc.), all the dimensionless coefficients in (5) are taken to be arbitrary since extra restrictions on the relative couplings may be enforced by those prospective studies. At this point, one shall notice that it also deserves to generalize WECG in (5) (as well as its higher dimensional extensions) with the ensuing Weyl’s extension of the dimensionally-extended Euler quadratic and cubic densities (namely, $\widehat{\chi}_{4}$ and $\widehat{\chi}_{6}$) associated with Weyl-invariant analogous of Gauss-Bonnet and Cubic Lovelock combinations [19, 20]
@@
$\qquad+3\widehat{R}_{abcd}\widehat{R}^{abcd}R+24\widehat{R}_{abcd}\widehat{R}^{ac}\widehat{R}^{bd}+16\widehat{R}_{a}{{}^{b}}\widehat{R}_{b}{{}^{c}}\widehat{R}_{c}{{}^{a}}-12\widehat{R}_{ab}\widehat{R}^{ab}\widehat{R}+\widehat{R}^{3},$ (9)
@@
where $\mathfrak{L}_{\mbox{ECG}}$ is the Lagrangian density of the strongly-coupled ECG [20] upgraded by the dimensionless couplings $\alpha_{i}$ which involves specific cubic curvature combinations only and is described as below
@@ IV.1 Metric Field Equations
during the variation of the action with respect to $g^{\mu\nu}$, one can show that, posterior to a long process of calculations, the metric field equation up to a boundary term can be written in the compact form according to the certain expressions of scalar field terms as follows
@@
Here, $\mathscr{O}_{i}=\mathscr{O}_{i}(R,A,\Phi)$ where $i=1\cdots 9$ in general. As sequentially presented below, each $\mathscr{O}_{i}$ term further contains the other kind of terms compactly denoted as $\Theta_{[i]j}A=\Theta^{[i]j}A$” where the first indices “ii” in the square bracket stands for specific $\mathscr{O}_{i}$-term in (14) it belongs to, and the second index “jj” signifies the number of the sub-term for same $\mathscr{O}_{i}$. Also, the pure specific curvature combinations in $\mathscr{O}_{i}$-terms are compactly denoted by $\widehat{\Omega}_{(i)}$ and presented in (B), explicitly. To be more precise, let us first notice that the term $\mathscr{O}^{(1)}_{\mu\nu}$ reads as
@@
Here, the corresponding term $\widehat{\Omega}^{\alpha\beta}_{(4)\mu\nu}$ is given in (43), whereas the coefficient functional $\Theta^{[1]1}_{\alpha}A_{\beta}$ and $\Theta_{\alpha}^{[4]i}A_{\beta}$ (running as $i=1\cdots 6$) are independently provided in (C.1) and (C.4). Similarly, the two-fold symmetric-anti symmetric index operation $\{\!\!\{\cdots\}\!\!\}$" is defined in (1).
@@
wherein the related term $\widehat{\Omega}^{(5)}_{\mu}$ comprising specific curvature interactions only is presented in (44). Here, the explicit forms of coupling $\kappa_{1({Q_{5}})}$ and coefficient functional $\Theta_{\alpha}^{[5]i}A_{\beta}$ (with $i=1\cdots 21)$ are respectively exhibited in (D.3) and (C.5).
@@
where the relevant bare curvature term $\widehat{\Omega}^{(7)}_{\mu\nu}$ and function $\Xi^{\alpha\beta}_{(5)}$ can be found in (46) and (12), manifestly. The constants $\zeta_{i}$’s, $\kappa_{3(5)}$ and $\kappa_{j(7)}$ (with $j=5,6,9$) are presented in (A), (D.2) and (D.5), respectively. Also, the different coefficient functional $\Theta_{[1]1}^{\alpha}A^{\beta}$, $\Theta_{[5]3}^{\alpha}A^{\beta}$, $\Theta^{\alpha}_{[6]k}A^{\beta}$ (for $k=2,4,10$) and $\Theta^{\alpha}_{[7]l}A^{\beta}$ (where $l=1\cdots 41$) are given in (C.1), (C.5), (C.6) and (C.7) in succession.
@@ IV.2 Field Equations for Weyl’s Vector and Scalar Fields
Firstly, due to the highly complex higher-order nature of the WECG as was underlined, we shall similarly collect separately the out-coming terms in the gauge field equations mainly according to $\nabla_{\mu}\Phi^{-2}$ and $\Phi^{-2}$ even though they are more simpler in contrast to the metric field equations. To be more precise, by varying the action in (10) with respect to $A^{\mu}$, one gets the gauge field equation in compact form as follows
@@
Here, the explicit forms of coefficient functional $\mathscr{X}_{[A]i}$ (where $i=1\cdots 11$) are given in (E).
As for the scalar field equation, it is actually straightforward to show that the variation of WECG action in (10) with respect to $\Phi$ turns out to be
@@
V Spontaneous Symmetry Breaking In the Maximally Symmetric Vacua
@@
$R_{\mu\nu\alpha\beta}=\frac{\Lambda}{3}(g_{\mu\alpha}g_{\nu\beta}-g_{\mu\beta}g_{\nu\alpha}),\qquad R_{\mu\nu}=\Lambda g_{\mu\nu},\qquad R=4\Lambda.$ (26)
Notice that here we are not searching for the exact solutions to the model which requires much more detailed separate studies, but instead we would like to determine whether the symmetry is broken in the classical vacua or not. In this part, we will see that the general field equations reduce to relatively simpler expressions around constant curvature backgrounds (or at least when the vacuum solutions are those of maximally symmetric spacetimes) although they apparently are highly intricate for generic backgrounds due to various higher-order non minimal interactions among the existing fields ($g_{\mu\nu},A_{\mu},\Phi$). In doing so, we follow [75]. To see that clearly, let us also pick $F_{\mu\nu}=0$ with the specific legitimate ansatz $A_{\mu}=0$ to avoid having any certain direction that would break the Lorentz-invariance of vacua and also freeze the scalar field to its vacuum expectation value as $\Phi=\left<\Phi_{vac}\right$$\sqrt{{\cal M}}$ in addition to (26). With these proper setups, the vector field equation (24) will be fulfilled immediately, while the metric and scalar field equations in (14) and (25) yield the identical vacuum field equation incorporating background cosmological constant and vacuum expectation value of the scalar field–which is also approval of the calculations–as below
@@
[…]
@@
[…]
@@
where $\widehat{\xi}$ and $r$ are given in (34). Observe that the two roots for both cases are imaginary so we will not study them in detail since they will provide non-hermitian complex operators associated with non-physical observable. (See conclusion for the relevant discussions and possible future directions.)
@@
[…]
Observe that the existence of the AdS solution solely comes from the unitarity and consistency conditions. Fundamentally, we are not seeking an exact solution in this work. Definitely, finding exact solutions for the WECG would be an interesting problem, but it is beyond our scope and needs a much more detailed separate study. Let us also note that there might be solutions that do not satisfy the unitary conditions. As mentioned, given its breadth, this could be the subject of another study. But, as a side comment and suggestion in finding exact solutions, since the field equations of WECG are quite complicated and difficult, one shall benefit from quite effective techniques. In the literature, there are various methods to obtain exact solutions. For example, since the field equations of nonlocal gravity theory are quite complex and difficult to find solutions as in our current case, using the methods in [87] for the exact solutions in nonlocal gravity or the so-called technique of Kerr-Schild form as in [88, 89], where the exact pp-wave and impulsive wave solutions of nonlocal infinite derivative gravity in [6] are found, will significantly simplify the field equations. Such a selection may simplify finding a solution significantly, but for our current purposes, it will not change the results obtained.
@@ VI Conclusions and Discussions
By upgrading the bare curvature terms to those composed of abelian gauge fields along with appropriately tuned weights of real scalar fields, we have constructed a Weyl-gauging of strongly coupled Einsteinian-Cubic gravity in Weyl’s geometry in $3+1$ dimensions, called Weyl-Einsteinian-Cubic gravity (WECG). This provides a cubic gauge theory of gravity. The model is invariant under the so-called Weyl’s local conformal transformations and hence does not involve any dimensionful parameter. Due to the nature of construction, the WECG describes $g_{\mu\nu},A_{\mu}$ and $\Phi$ as the fundamental dynamical fields. We have also obtained the most general field equations associated with the existing propagating fields about any generic spacetime background. Accordingly, we have demonstrated that the ordinary ECG emerges as the lower energy limit of WECG in its local non-conformal-invariant vacua about (anti-) de Sitter and flat spacetimes where the vacuum expectation value is fixed to the Planck mass scale. Additionally, we have gotten the same cubic vacuum field equations in the maximally symmetric backgrounds (namely, (anti-) de Sitter and flat spaces) from the generic field equations for the metric and real scalar field where that of the gauge field is automatically fulfilled as expected. Ensuingly, we have explicitly obtained the desired real cosmological constant and vacuum expectation value of the scalar field and also determined the reality conditions on the values of existing relative couplings in WECG for the (anti-) de Sitter vacuum spacetime. Here, the natural existence of (anti)-de Sitter space spontaneously breaks Weyl’s local conformal symmetry which fixes all the dimensionful parameters. Moreover, we have observed that anti-de Sitter space is allowed to be a vacuum background solution of WECG, while de Sitter space is ruled out. As for the flat spaces, the local conformal invariance of the vacuum persists to be unbroken. Here, we have demonstrated that the essential dimensionful parameter arises at the renormalization scale at the one-loop level, and thus the symmetry is broken radiatively.
@@
Furthermore, note that the section V is not devoted to find the exact solutions of the model, but rather the existing symmetry-breaking mechanism and unitarity analysis have been analyzed to some level here to solely determine how the local Weyl’s conformal symmetry is broken and accordingly which constant curvature vacua are allowed or not, as in, e.g., [75]. But, of course, as is well-known from literature, determining explicit unitary (ghost and tachyon-free) parameter regions and also the emerging symmetry-breaking mechanism require separate elaborated studies as in, e.g., [71, 111]. Also, finding exact solutions and spacetime singularities of the model seems to also be essential. Since the field equations of WECG are quite complex and cumbersome, one can utilize the effective techniques used in nonlocal gravity theories as in [87] or the Kerr-Schild form as in, e. g., [88, 89]. As was mentioned in the beginning, as another possible future direction, it will also be very interesting to study the plausible links between the current cosmological WECG and foremost alternative higher-order (non-local) conformal (quantum) gravity (in different dimensions) rooted in the Weyl tensors and traceless Ricci tensor as in [81, 82, 83, 84]. Finally, as is conducted in [83, 84], it will also be interesting to analyze the model at the loop-level and figure out the modifications to the propagator structures. Since all the characteristics of a model cannot be throughly studied just in a single paper, we suggest them as the potentially interesting future projects about the model.
Acknowledgments
This article is dedicated to Roman JACKIW’s second death anniversary who was one of the world-leading scientists of the century and my former postdoc advisor at MIT/CTP. We thank Daniel Grumiller, Mithat Unsal, Daniel Kabat, Anupam Mazumdar, Ercan Kilicarslan, Heeseung Zoe and Paul Brown for useful discussions and suggestions. We would also like to thank an anonymous referee for her/his constructive remarks, critics and suggestions on the paper.